OpenStudy (anonymous):

) What is the slope of the line through the points (5, 2) and (5, -3)? b) Describe the line through these points.

OpenStudy (anonymous):

calculating the equation of a line with 2 points is quite simple what u do is assign x1,y1 and x2,y2 for ur points. lets say 5,2 is x1,y1 and 5,-3 is x2and y2. its y-y1=(y2-y1)/(x2-x1) *x-x1. so it would be y-2=(-5)/(0) *( x-5) but since gradient is 0 the equation is just y=2

OpenStudy (anonymous):

oh btw slope is the gradient which is the y2-y1)/x2-x1 so -5 over 0 which is actually undefined

OpenStudy (anonymous):

yeah that's what I got undefined

OpenStudy (anonymous):

what would be part B?

OpenStudy (anonymous):

@shamil98 can you help me?

OpenStudy (shamil98):

the slope of a vertical line is undefined. so the line is vertical :)

OpenStudy (anonymous):

:) thank you

OpenStudy (anonymous):

is the first part undefined or 2 ?

OpenStudy (anonymous):

@shamil98

OpenStudy (shamil98):

the slope is undefined.

OpenStudy (anonymous):

okay that's what I thought it was, thank you for your help :)